# Signed Angle Rigid Graphs for Network Localization and Formation Control

### v1

@@ VIII-D Proof of Theorem 3
Proof: (i) $\Leftrightarrow$ (iii). Suppose $(\mathcal{G},p)$ is ISAR. From Lemma 4 and $\mathscr{E}(p)\subseteq\mathscr{E}_{\mathcal{G}}(p)$, it suffices to show that for any $q\in\mathscr{E}_{\mathcal{G}}(p)$, it always holds $q\in\mathscr{E}(p)$. Since $\mathcal{G}$ must be connected, it follows from Lemma 3 that $q\in\mathscr{E}_{\mathcal{G}}(p)$ if only if $(I_{n}\otimes\mathscr{R}_{o}(\theta))^{-1}q\in B^{-1}_{\mathcal{G}}(B_{\mathcal{G}}(p))$ for some $\theta\in[0,2\pi)$. According to [9, Theorem 6], we know that for a given IBR framework $(\mathcal{G},p)$, one has
@@
Suppose the shape of $(\mathcal{G},p)$ can be uniquely determined by signed angles up to uniform rotations, translations, and scalings. Then, $p$ must be non-degenerate, and one has $\mathscr{E}(p)=\mathscr{E}_{\mathcal{G}}(p)$ according to Lemma 4. Therefore, both $\mathscr{E}(p)$ and $\mathscr{E}_{\mathcal{G}}(p)$ are 4-dimentional smooth manifolds. According to [33, Proposition 3.10], the tangent space of $\mathscr{E}_{\mathcal{G}}(p)$ at $p$ is 4-dimensional, which indicates that $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$. As a result, it follows from Lemma 2 that $(\mathcal{G},p)$ is ISAR.
(ii) $\Leftrightarrow$ (iii). Suppose $(\mathcal{G},p)$ is non-degenerate and GSAR. It suffices to show $\mathscr{E}(p)=\mathscr{E}_{\mathcal{G}}(p)$. Since $\mathscr{E}_{\mathcal{G}}(p)=\mathscr{E}_{\mathcal{K}}(p)$ and (i) $\Leftrightarrow$ (iii), we just need to show that there exists an ISAR spanning subframework of $(\mathcal{K},p)$. From the non-degeneracy of $(\mathcal{K},p)$, without loss of generality, we can select three non-collinear vertices from $(\mathcal{K},p)$ and index them by $1,2,3$. It can be observed from Lemma 2 that the framework $(\mathcal{G}^{\prime},p^{\prime})$ embedded by the triangle $\mathcal{G}^{\prime}$ composed of $1,2,3$ is ISAR, where $p^{\prime}\in\mathbb{R}^{6}$ is the configuration of $1,2,3$ in $p$. Since for any $4\leq j\leq n$, there always exist $i,k\in\{1,2,3\}$ such that $b_{ji}$ and $b_{jk}$ are not collinear. Combining [10, Theorem 3], Theorem 1, and [17, Lemma 7], we conclude that the framework embedded by the graph constructed by adding vertex $j$ and edges $(j,i)$, $(j,k)$ to $\mathcal{G}^{\prime}$ for $j=4,\cdots n$ iteratively is ISAR.
@@ VIII-E Proof of Theorem 10
To show convergence, similar to [40, pp. 110], for the network $(\mathcal{G},p,\mathcal{A})$ and the AIS $\bar{\mathcal{T}}_{\mathcal{G}}$, we define $\hat{\mathcal{G}}\triangleq(\hat{\mathcal{V}},\hat{\mathcal{E}})$ as the bearing graph under (31), which is a directed graph describing the pairwise relationship between bearings in the angle errors feedback according to (31). Here $v_{ij}\in\hat{\mathcal{V}}$ indexes the estimate $\hat{b}_{ij}(t)$ for $i\in\mathcal{V}$ and $j\in\mathcal{N}^{i}_{\mathcal{G}}$ and is computed by
@@
Note that according to Theorem 1 and [15, Corollary 3], Assumption 1 implies that $(\mathcal{G},p,\mathcal{A})$ is bearing localizable [15]. Therefore, we first prove that $\hat{b}_{ij}(t)$ exponentially converges to the global bearing $b_{ij}$ for all $j\in\mathcal{N}_{i}$. Then, the proof of the convergence of $\hat{p}_{i}(t)$ to $p_{i}$ for all $i\in\mathcal{F}$ is similar to that in [38, Theorem 2].
@@
Since for each vertex $v_{ij}\in\hat{\mathcal{V}}\backslash\hat{\mathcal{A}}$, there exists a directed path from some vertex in $\hat{\mathcal{A}}$ to $v_{ij}$ by Lemma 10, and $x_{jk}=B^{\top}_{jk}\hat{b}_{jk}=B^{\top}_{jk}b_{jk}=[1,0]^{\top}$ for $v_{jk}\in\hat{\mathcal{A}}$ and all $t\geq 0$, $x_{ij}$ converges to $[1,0]^{\top}$ for all $j\in\mathcal{N}_{i}$ according to the result in [39, Theorem 5.1]. It implies that the system converges to
@@
where $\theta$ is the signed angle from $\hat{b}_{ij}$ to $b_{ij}$. It follows from (53) that $\|\hat{b}_{ij}\|\mathscr{R}_{o}(\theta)[1,0]^{\top}=[1,0]^{\top}$. Therefore, one has $\|\hat{b}_{ij}\|=1$ and $\theta=0$. As a result, $\hat{b}_{ij}\rightarrow b_{ij}$ as $t\rightarrow\infty$ for all $j\in\mathcal{N}_{i}$. Since the system (52) is linear, the convergence is exponentially fast [40, Corollary 4.3]. ∎

### v2

@@ VIII-D Proof of Theorem 3
Proof: (i) $\Leftrightarrow$ (iii). Suppose $(\mathcal{G},p)$ is ISAR. From Lemma 4 and $\mathscr{E}(p)\subseteq\mathscr{E}_{\mathcal{G}}(p)$, it suffices to show that for any $q\in\mathscr{E}_{\mathcal{G}}(p)$, it always holds $q\in\mathscr{E}(p)$. Since $\mathcal{G}$ must be connected, it follows from Lemma 3 that $q\in\mathscr{E}_{\mathcal{G}}(p)$ if only if $(I_{n}\otimes\mathscr{R}_{o}(\theta))^{-1}q\in B^{-1}_{\mathcal{G}}(B_{\mathcal{G}}(p))$ for some $\theta\in[0,2\pi)$. According to [10, Theorem 6], we know that for a given IBR framework $(\mathcal{G},p)$, one has
@@
Suppose the shape of $(\mathcal{G},p)$ can be uniquely determined by signed angles up to uniform rotations, translations, and scalings. Then, $p$ must be non-degenerate, and one has $\mathscr{E}(p)=\mathscr{E}_{\mathcal{G}}(p)$ according to Lemma 4. Therefore, both $\mathscr{E}(p)$ and $\mathscr{E}_{\mathcal{G}}(p)$ are 4-dimentional smooth manifolds. According to [36, Proposition 3.10], the tangent space of $\mathscr{E}_{\mathcal{G}}(p)$ at $p$ is 4-dimensional, which indicates that $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$. As a result, it follows from Lemma 2 that $(\mathcal{G},p)$ is ISAR.
(ii) $\Leftrightarrow$ (iii). Suppose $(\mathcal{G},p)$ is non-degenerate and GSAR. It suffices to show $\mathscr{E}(p)=\mathscr{E}_{\mathcal{G}}(p)$. Since $\mathscr{E}_{\mathcal{G}}(p)=\mathscr{E}_{\mathcal{K}}(p)$ and (i) $\Leftrightarrow$ (iii), we just need to show that there exists an ISAR spanning subframework of $(\mathcal{K},p)$. From the non-degeneracy of $(\mathcal{K},p)$, without loss of generality, we can select three non-collinear vertices from $(\mathcal{K},p)$ and index them by $1,2,3$. It can be observed from Lemma 2 that the framework $(\mathcal{G}^{\prime},p^{\prime})$ embedded by the triangle $\mathcal{G}^{\prime}$ composed of $1,2,3$ is ISAR, where $p^{\prime}\in\mathbb{R}^{6}$ is the configuration of $1,2,3$ in $p$. Since for any $4\leq j\leq n$, there always exist $i,k\in\{1,2,3\}$ such that $b_{ji}$ and $b_{jk}$ are not collinear. Combining [11, Theorem 3], Theorem 1, and [17, Lemma 7], we conclude that the framework embedded by the graph constructed by adding vertex $j$ and edges $(j,i)$, $(j,k)$ to $\mathcal{G}^{\prime}$ for $j=4,\cdots n$ iteratively is ISAR.
@@ VIII-E Proof of Theorem 10
To show convergence, similar to [41, pp. 110], for the network $(\mathcal{G},p,\mathcal{A})$ and the AIS $\bar{\mathcal{T}}_{\mathcal{G}}$, we define $\hat{\mathcal{G}}\triangleq(\hat{\mathcal{V}},\hat{\mathcal{E}})$ as the bearing graph under (31), which is a directed graph describing the pairwise relationship between bearings in the angle errors feedback according to (31). Here $v_{ij}\in\hat{\mathcal{V}}$ indexes the estimate $\hat{b}_{ij}(t)$ for $i\in\mathcal{V}$ and $j\in\mathcal{N}^{i}_{\mathcal{G}}$ and is computed by
@@
Note that according to Theorem 1 and [16, Corollary 3], Assumption 1 implies that $(\mathcal{G},p,\mathcal{A})$ is bearing localizable [16]. Therefore, we first prove that $\hat{b}_{ij}(t)$ exponentially converges to the global bearing $b_{ij}$ for all $j\in\mathcal{N}_{i}$. Then, the proof of the convergence of $\hat{p}_{i}(t)$ to $p_{i}$ for all $i\in\mathcal{F}$ is similar to that in [39, Theorem 2].
@@
Since for each vertex $v_{ij}\in\hat{\mathcal{V}}\backslash\hat{\mathcal{A}}$, there exists a directed path from some vertex in $\hat{\mathcal{A}}$ to $v_{ij}$ by Lemma 10, and $x_{jk}=B^{\top}_{jk}\hat{b}_{jk}=B^{\top}_{jk}b_{jk}=[1,0]^{\top}$ for $v_{jk}\in\hat{\mathcal{A}}$ and all $t\geq 0$, $x_{ij}$ converges to $[1,0]^{\top}$ for all $j\in\mathcal{N}_{i}$ according to the result in [40, Theorem 5.1]. It implies that the system converges to
@@
where $\theta$ is the signed angle from $\hat{b}_{ij}$ to $b_{ij}$. It follows from (53) that $\|\hat{b}_{ij}\|\mathscr{R}_{o}(\theta)[1,0]^{\top}=[1,0]^{\top}$. Therefore, one has $\|\hat{b}_{ij}\|=1$ and $\theta=0$. As a result, $\hat{b}_{ij}\rightarrow b_{ij}$ as $t\rightarrow\infty$ for all $j\in\mathcal{N}_{i}$. Since the system (52) is linear, the convergence is exponentially fast [41, Corollary 4.3]. ∎